# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ A.1 Developments of the Cities
Note: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya. Figure A.1b shows the share of people in Tokyo (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan. Source: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
@@
Panel A: Population and Industrial Structures
@@ A.2 Financial Institutions in the Cities
A representative savings institution for households in prewar Japan is the postal savings (yūbin chokin). Although postal savings was initially designed as a small savings institution for low-income households, the share of postal savings to Japan’s total savings was roughly $10$% in the 1910s (Tanaka 2018, p. 29). The number of postal savings accounts per total population in the early 1910s exceeded $20$% in Japan (Tanaka 2018, p. 33), and the saving rate in prewar Japan was indeed higher than that of the U.S. (James and Suto 2011). As shown in Panel A of Table A.2, roughly $1.4$ million people had accounts in Tokyo in 1920, covering approximately $67$% of the entire population in Tokyo at that time. The average savings amount per capita was $64.5$ yen. The average figure for the postal savings accounts in Kyobashi Ward was $44.5$ yen at the end of 1919 (Tokyo City Office 1921, pp. 892–893). This amount is less than the monthly earnings of the skilled-factory workers (Section 3.2). Unfortunately, the statistics on postal savings by depositors’ occupation in Tokyo City are unavailable. [^69] The statistics on Japan show that the share of depositors in the manufacturing sector was $4.8$% in 1920. The share based on the amount of the deposit is $5.5$%. See Postal Savings Bureau (1924, p. 72). However, this figure suggests that the average amount of postal savings was insufficient to smooth factory workers’ monthly consumption.
@@
Savings bank (chochiku ginkō) is another famous savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81). Panel A of Table A.2 shows that the number of people having savings accounts was roughly $1.4$ million people had the accounts, accounting for approximately $58$% of total people in Tokyo at that time. Although this figure is similar to the postal savings, the workers in the manufacturing industry were less likely to rely on the savings banks than the workers in the other industrial sectors. Table A.3 summarizes the number of depositors and the amount of deposits by occupation in Tokyo City. In Kyobashi, the commerce and miscellaneous sectors shared roughly $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%. This is mainly unchanged if I use the amounts of deposits to calculate the shares. Tokyo Institute for Municipal Research stated that this tendency captures the fact that savings banks are mainly designed as financial institutions for the lower-income class (Tokyo Institute for Municipal Research 1925b, p. 81). This suggests that skilled factory workers were less likely to use this savings institution. This may be consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices. [^70] There were 123 savings banks in Tokyo City in 1922, and 10 of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p. 76). However, Tsukishima, a representative manufacturing area, did not have savings banks at that time.
@@
Mutual loan association (mujin) is another type of savings institution. The members of this association pay a fixed amount of money several times, and each member withdraws the total amount of payments each time under specific rules. The lotteries or bidding system decided which members could withdraw money each time. This means that the workers were less likely to use this institution to deal with the acute negative income shocks. [^71] This is the most representative mutual loan system called tsūjyō (i.e., normal) mujin. There was a different type called sueoki chokin (or tsumitate kai) that people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp. 3–4). However, this is also not flexible for compensating the income shocks. Instead, it had been used to make a certain amount of money among small business owners and merchants. A report suggests that the workers in the manufacturing sector were roughly 10% of all the mutual loan associations in Tokyo (Tokyo City Office 1935, p. 25). Tokyo Chamber of Commerce stated that the mutual loan association was used for business and rarely used for household consumption (Tokyo Chamber of Commerce 1918, p. 121).
@@
Cooperative societies were also available for worker households. Cooperative societies in Tokyo used credit as their payment method: households could buy goods in advance and pay at the end of the month or twice (middle and the end of) a month. The commodity prices were similar to or slightly lower than those in the retail shops (Central Federation of Industrial Associations 1925, pp. 18–19). However, the cooperatives only covered a few citizens. There were 26 cooperative societies in Tokyo prefecture in 1924, and 23 of them were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). Among these 23 cooperatives, while the two largest societies had roughly 4000 members in total, the others were generally small societies with less than 100 members. This figure accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand (Table A.1). Most of the THBS households had not used the cooperatives.
@@
An alternative institution is the credit purchase in retail shops. The THBS households had often used the credit purchase in various retail shops for rice, fish, vegetables, alcohol, firewood, and charcoal. The details of the credit purchase are not described in the documents, and the systematic statistics on the credit purchase are also unavailable. This may be mainly because these transactions were conducted between each shop and worker. However, at least, the THBS budget data suggest that credit purchases had been widely used among the skilled-factory worker households at that time.
@@
Pawnshops were the most popular lending institution among low-income working-class households. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. Among $88$ pawnshops in Kyobashi, the average redemption rate was approximately $94$% ($401,222$ cases/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp. 888–889). The accessibilities were higher than the other lending institutions. The number of pawnshops in Tokyo was $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. As described in the main text, there were 13 pawnshops and only five moneylenders in Tsukishima. Thus, the accessibility to a pawnshop was undoubtedly high for the workers in the city.
@@
However, the inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was $14$% of the monthly earnings of the THBS households. While the pawnshops were the popular lending institution, they could not offer a sufficient amount of temporary income for relatively large idiosyncratic shocks. According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the share of agriculture, commerce, and manufacturing sectors were 1.7, 19.5, and 15.6%, respectively. The most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p. 25). Tokyo City Office stated that the pawnshops were mainly used among low-income (poor) working-class households (Tokyo City Social Affairs Bureau 1921, p. 9). This indicates the evidence that the primary pawnshop users were not the skilled-worker households like the THBS households.
@@
Another lending institution is the money lenders (kinsen kashitsuke gyō). Generally, business owners use money lenders, and the interest rates in the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp. 184; 248). Tokyo Institute for Municipal Research was concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp. 106–108). Given this, while there were two money lenders in Tsukishima, they might have been less likely used by the factory worker households.
@@
Entire Tokyo City
@@
Notes: This table summarizes the number of depositors and the amount of deposits in the savings banks in Tokyo City by occupation on 31st December 1919. Source: Tokyo City Office (1922a, pp. 864–865).
@@
Notes: This map shows the location of the pawnshops and money lenders in Tsukishima. Bullseye indicates the pawnshop (shichiya). The black circle shows the money lenders (kanekashi gyō). The white circle shows the rickshaw station (kuruma yado). Source: Department of Health, Ministry of the Interior 1923c, tenth map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
@@ A.3 Development of Tsukishima
Tsukishima (transliterated as “moon island”) is a reclaimed land on a sandbar of the Sumida River. Ishikawa Island and Tsukuda Island at the northern end of Tsukishima are the starting point of its history. The following description of this section is based on Chūō Ward (1994, pp. 122–124).
Ishikawa Island was built in 1626 by a boatman, Hachizaemon Ishikawa, who received a fief from the Shogunate. In 1790, the Ishikawajima Prison was built on Ishikawa Island. This was a facility for the reintegration of minor offenders and those without lodgings and was called a Ninsoku Yoseba. After Perry arrived in Japan in 1853, the Shogunate ordered the Mito Clan to build a shipyard on Ishikawa Island. A month later, the groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856. In 1866, the Ishikawajima Lighthouse was built on the west bank of the Yoseba and served to signal the position of ships entering the Nihonbashi direction. In 1870, the Ninsoku Yoseba became a place of imprisonment (Toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877. The prison was finally moved to Sugamo and the penitentiary on Ishikawa Island was closed in 1895.
Adjacent to Ishikawa Island is called Tsukuda Island, where fishermen from “Tsukuda” in Osaka Prefecture settled in 1644 after receiving land from the shogunate (the island was named after their hometown, Tsukuda). In 1645, the “Tsukuda ferry” began, and throughout the Edo period, Tsukuda Island was known as a fishing island for whitefish for the shogun’s household. In 1863, the Tsukudajima Battery was built on the south side of Tsukuda Island, one of 13 batteries used to defend Edo Bay. This battery later became the starting point for the reclamation of Lot 1 of Tsukishima.
In 1884, the Port of Tokyo Waterway Dredging Project was initiated. This was to secure the passage of ships by sweeping away the sediment in the shipping channel, and the Tsukishima landfill was created using the deposit. The first Tsukishima site was completed in 1891, and the “Tsukishima ferry” began the same year. The second site was created in 1894; then the Shin-(new) tsukudajima was built in 1896. From the west shore of Shin tsukudajima, the island overlooks the Boso Peninsula in Chiba Prefecture, making it a favorite vacation spot for writers. Cultural figures such as Toson Shimazaki and Kaoru Osanai stayed here from 1907 to the early Taisho period. In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to industry development (until then, Suzukiiya had been selling water to supply the entire island using a ferryboat). The “Kachidoki Ferry” began to lead to Lot 2 in 1905. In 1913, Lot 3 was finally completed.
@@ A.4 Housings in Tsukishima
Since the Tsukishima Survey included a housing survey (Table B.1), the type of housing in the entire island can be determined. Table A.4 summarizes the results of this survey. The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp. 69–76).
$88$ ($46$%) of the detached homes are in Tsukudajima. Tsukudajima is reported to be “superior to other areas in terms of housing quality, and the percentage of inferior housing is meager”. Conversely, Shin-tsukudajima has a low percentage of detached homes, is actually “inhabited by unskilled laborers” and “occupies the most inconvenient housing status of all the islands.” Lot 1 is a cluster of numerous residences, mainly tenements, which is consistent with the fact that Lot 1 functions as a center of life in Tsukishima. Lot 2 had a relatively high percentage of single-family and two-family houses and appears to have had a character as a “workers’ residential area” than Lot 1. These indicate that there may have been some geographic differences in income distribution regarding housing type.
Overall, however, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) are row houses, and only $190$ are single-family homes. This confirms that Tsukishima was an industrial area with many working families. The workers lived in very similar tenements and with similar standards of living. Figure A.4 is a photo showing these tenement houses.

### 25 Feb 2024 v6

@@ A.1 Developments of the Cities
Note: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities. Figure A.1b shows the share of people in Tokyo (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan. Sources: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
@@
Panel A: Population and Workers
@@
Figure A.1a illustrates the developments of the large cities in Japan between 1889 and 1920. Table A.1 summarizes the scales, industrial structures, infrastructures, and meteorological features of the large cities.
@@ A.2 Financial Institutions in the Cities
A representative savings institution is the postal savings (yūbin chokin). Postal savings was initially designed as a small savings institution for low-income households, and the share of postal savings to total savings was roughly $10$% in the 1910s (Tanaka 2018, p. 29). However, the number of postal savings accounts per total population in the early 1910s exceeded $20$% (Tanaka 2018, p. 33). [^73] James and Suto (2011) show that the saving rate in prewar Japan was higher than that of the U.S. As shown in Panel A of Table A.2, roughly $1.4$ million people had accounts in Tokyo in 1920. This covers approximately $67$% of the entire population in Tokyo and the average savings amount per capita was $64.5$ yen. In Kyobashi Ward, the average figure for the postal savings accounts was $44.5$ yen at the end of 1919 (Tokyo City Office 1921, pp. 892–893). This amount is less than the average monthly earnings of the skilled-factory workers (Section 3.2). Unfortunately, the statistics on postal savings by depositors’ occupation are unavailable. [^74] There are some statistics for the entire Japan. According to the Postal Savings Bureau (1924, p. 72), the share of depositors in the manufacturing sector was $4.8$% (the share based on the deposit amount is $5.5$%). However, this implies that, on average, postal savings may not provide sufficient temporary income.
@@
Savings bank (chochiku ginkō) is another famous savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81). Panel A of Table A.2 shows that the number of people having savings accounts was roughly $1.46$ million, accounting for approximately $67.3$% of total people in Tokyo at that time. Although this figure is similar to the postal savings, the workers in the manufacturing industry were less likely to rely on the savings banks than the workers in the other industrial sectors. Table A.3 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B). In Kyobashi, the commerce and miscellaneous sectors shared roughly $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%. This is unchanged if I use the amounts of deposits to calculate the shares. Tokyo Institute for Municipal Research stated that this tendency captures the fact that savings banks are mainly designed as financial institutions for low-income households (Tokyo Institute for Municipal Research 1925b, p. 81). Thus, skilled factory worker households, like the THBS households, were less likely to use the savings banks. This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices. [^75] There were 123 savings banks in Tokyo City in 1922, and 10 of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p. 76). However, Tsukishima, a representative manufacturing area, did not have savings banks at that time.
@@
Mutual loan association (mujin) is another type of savings institution. The members of this association pay a fixed amount of money several times, and each member withdraws the total amount of payments each time under specific rules. The lotteries or bidding system decided which members could withdraw money each time. This means that the workers could not use this institution to deal with the acute negative income shocks. [^76] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin. There was a different type called sueoki chokin (or tsumitate kai) that people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp. 3–4). Clearly, both are also not flexible for compensating the income shocks. Instead, it had been used to make a certain amount of money among small business owners and merchants. A report suggests that the workers in the manufacturing sector shared only $10$% of all the mutual loan associations in Tokyo (Tokyo City Office 1935, p. 25). Moreover, the Tokyo Chamber of Commerce stated that the mutual loan association was used for business and rarely for household consumption (Tokyo Chamber of Commerce 1918, p. 121).
@@
Cooperative societies were also available for worker households. Cooperative societies in Tokyo used credit as their payment method: households could buy goods in advance and pay at the end of the month or twice (middle and the end of) a month. The commodity prices were similar to or slightly lower than those in the retail shops (Central Federation of Industrial Associations 1925, pp. 18–19). However, the cooperatives only covered a few citizens. There were 26 cooperative societies in Tokyo prefecture in 1924, and 23 of them were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). Among these 23 cooperatives, two were large societies with roughly 4000 members, and the others were small societies with less than 100 members. The total number of accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand (Table A.1). In fact, most of the THBS households had not used the cooperatives.
@@
An alternative borrowing institution is the credit purchase in retail shops. The THBS households had frequently used credit purchases in various retail shops for rice, fish, vegetables, alcohol, firewood, charcoal, and medicines. The details of the credit purchase are little documented, and the systematic statistics on the credit purchase are also unavailable. This must be because the information on the transactions is personal. Despite this, an available document shows that credit purchases are used in various retail shops (Section 5). The THBS budget data show a consistent tendency for credit purchases to be widely used among factory worker households.
@@
Pawnshops were the popular lending institution among low-income working-class households. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. Among $88$ pawnshops in Kyobashi, the average redemption rate was approximately $94$% ($401,222$ cases/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp. 888–889). The accessibilities were higher than the other lending institutions. The number of pawnshops in Tokyo was $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. This means that the accessibility to a pawnshop was relatively high for the workers. As described in the main text, the number of pawnshops in Tsukishima was 13, whereas that of moneylenders was five.
@@
However, the inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was $12$% of the monthly earnings of the THBS households. While the pawnshops were the popular lending institutions, they could not offer a sufficient amount of temporary income for relatively large idiosyncratic shocks. According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the share of users in the agriculture, commerce, and manufacturing sectors was 1.7%, 19.5%, and 15.6%, respectively. In turn, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p. 25). In fact, the Tokyo City Office stated that the pawnshops were mainly used among low-income (poor) working-class households (Tokyo City Social Affairs Bureau 1921, p. 9). This indicates the evidence that the primary pawnshop users were not the skilled-worker households like the THBS households.
@@
Another lending institution is the money lenders (kinsen kashitsuke gyō). Generally, business owners used the money lenders. The interest rates in the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp. 184; 248). Tokyo Institute for Municipal Research was concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp. 106–108). Given this fact, while there were two money lenders in Tsukishima, they might have been less likely used by the factory worker households.
@@
Pane A: Entire Tokyo City
@@
Notes: This table summarizes the number of depositors and the amount of deposits in the savings banks by occupation. These statistics were measured on 31st December 1919. Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively. Source: Tokyo City Office (1922a, pp. 864–865).
@@
Notes: This map shows the location of the pawnshops and moneylenders in Tsukishima. The bullseye indicates the pawnshop (shichiya). The black circle shows the money lenders (kanekashi gyō). The white circle shows the rickshaw station (kuruma yado). Source: Department of Health, Ministry of the Interior 1923c, tenth map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
@@ A.3 Development of Tsukishima
Tsukishima (transliterated as “moon island”) is a reclaimed land on a sandbar of the Sumida River. Ishikawajima and Tsukudajima (“jima” means island) at the northern end of Tsukishima are the starting point of its history (Figure A.3 for map). The following description of this section is based on Chūō Ward (1994, pp. 122–124).
Ishikawajima was built in 1626 by a boatman, Hachizaemon Ishikawa, who received a fief from the Shogunate. In 1790, the Ishikawajima Prison was built on Ishikawajima. This was a facility for the reintegration of minor offenders and those without lodgings and was called a Ninsoku Yoseba. After Matthew Perry arrived in Japan in 1853, the Shogunate ordered the Mito Clan to build a shipyard on Ishikawajima. A month later, the groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856. In 1866, the Ishikawajima Lighthouse was built on the west bank of the Yoseba and served to signal the position of ships entering the Nihonbashi direction. In 1870, the Ninsoku Yoseba became a place of imprisonment (Toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877. The prison was finally moved to Sugamo in Kitatoshima County in Tokyo prefecture, and the penitentiary on Ishikawajima was closed in 1895.
Adjacent to Ishikawajima is called Tsukudajima, where fishermen from “Tsukuda” in Osaka prefecture settled in 1644 after receiving land from the Shogunate (i.e., the island was named after their hometown, Tsukuda). In 1645, the “Tsukuda ferry” began, and throughout the Edo period, Tsukudajima was known as a fishing island for whitefish for the Shogun’s household. In 1863, the Tsukudajima Battery was built on the south side of Tsukudajima, one of 13 batteries used to defend Edo Bay. This battery later became the starting point for the reclamation of Lot 1 of Tsukishima.
In 1884, the Port of Tokyo Waterway Dredging Project was initiated. This was to secure the passage of ships by sweeping away the sediment in the shipping channel, and the Tsukishima landfill was created using the deposit. The first Tsukishima site was completed in 1891, and the “Tsukishima ferry” began the same year. The second site was created in 1894; the Shin (new) tsukudajima was built in 1896. From the west shore of Shin tsukudajima, the island overlooks the Boso Peninsula in Chiba Prefecture, making it a favorite vacation spot for writers. Cultural figures such as Toson Shimazaki and Kaoru Osanai stayed here from 1907 to the early Taisho period. In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to industry development. [^77] Until then, Suzukiiya had been selling water to supply the entire island using a ferryboat. The “Kachidoki Ferry” began to lead to Lot 2 in 1905. In 1913, Lot 3 was finally completed.
@@ A.4 Housings in Tsukishima
Since the Tsukishima Survey included a housing survey (Table B.1), the housing types in the entire island can be determined. Table A.4 summarizes the results of this survey. The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp. 69–76).
$88$ ($46$%) of the detached homes are in Tsukudajima. Tsukudajima is reported to be “superior to other areas in terms of housing quality, and the percentage of inferior housing is meager”. Shin-tsukudajima had a low percentage of detached homes. The report states that it was indeed “inhabited by unskilled laborers” and “occupies the most inconvenient housing status of all the islands.” Lot 1 was a center of life in Tsukishima and is a cluster of numerous residences, mainly tenements. Lot 2 had a relatively high percentage of single-family and two-family houses and appears to have had a character as a “workers’ residential area” than Lot 1. These indicate that there might have been some geographic differences in income distribution regarding housing type. Overall, however, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) are row houses, and only $190$ are single-family homes. This confirms that Tsukishima was an industrial area with many working families. The workers lived in similar tenements and with similar standards of living. Figure A.4 is a photo showing these tenement houses.